# A Prefrontal Cortex-inspired Architecture for Planning in Large Language Models

### v3

@@ 6 Conclusion and Future Directions
In this work, we have proposed the LLM-PFC architecture, an approach aimed at improving the planning ability of LLMs by taking inspiration from the modular architecture of the human PFC. In experiments on three challenging planning domains, we found that LLM-PFC significantly improved planning performance over standard LLM methods. While these results represent a significant step forward, there is still room for improvement. In particular, the model has less than optimal performance on Tower of Hanoi, the Reward Revaluation graph traversal task, and the Logistics planning task (Valmeekam et al., 2023) (see Section A.5). This may be due in part to the inherent limitations of prompting and in-context learning as methods for the specialization of LLM-PFC’s modules. A promising avenue for further improvement may be to jointly fine-tune the modules across a range of diverse tasks (which requires open-source models), rather than relying only on black box methods (our only option with GPT-4). A white-box approach would also eliminate the need for task-specific prompts, and potentially enable zero-shot planning on novel tasks.
@@
An additional limitation of the current implementation is the computational cost incurred by the model (see Section A.7). Although this aligns well with the deliberative nature of controlled, PFC-dependent (i.e., ‘system 2’) processes (Kahneman, 2011), it would nevertheless be desirable to find ways to reduce these costs. One way to address this might be to use smaller models to perform some of the easier functions, or to fine-tune smaller models to perform the specialized roles of each module. Additionally, it may be possible to improve the efficiency of the model by caching, or pre-computing, the outputs of some modules.
LLM-PFC also has important implications for neuroscientific models of PFC function. Though much work has characterized the function of individual PFC subregions, there has been less emphasis on the development of integrative models in which these functions interact to carry out coordinated behavior. The present work represents a first step in that direction. An important next step will be to directly evaluate LLM-PFC as a model of neural data, which may then lead to further refinements of the model. We look forward to investigating these possibilities in future work.
@@ Appendix A Appendix
A.1 Extended Related Work
@@
In this section, we consider in more detail how LLM-PFC relates to existing black-box LLM approaches:
@@
  • Similar to LLM-PFC, both scratchpad (Nye et al., 2021) and chain-of-thought (CoT) (Wei et al., 2022b) decompose a problem into intermediate computations. However, unlike LLM-PFC, neither scratchpad nor CoT factorize these intermediate computations into specialized modules.
@@
  • Tree-of-thought (ToT) (Yao et al., 2023) introduces some degree of factorization, but the factorization is not as extensive as in LLM-PFC. The ‘generator’ module in ToT carries out a combination of the functions carried out by both the $\operatorname{Actor}$ (action proposal) and the $\operatorname{Predictor}$ (prediction of the states that will result from these actions) in LLM-PFC. The ‘evaluator’ module in ToT carries out a combination of the functions carried out by both the $\operatorname{Monitor}$ (error detection) and the $\operatorname{Evaluator}$ (prediction of state value) in LLM-PFC. ToT does not contain any component that carries out the functions of the $\operatorname{TaskDecomposer}$ (subgoal proposal) and the $\operatorname{Orchestrator}$ (autonomously determining when a goal or subgoal has been achieved).
  • Multi-agent debate (i.e., Society of Mind) (Du et al., 2023) involves the interaction of multiple LLM instances; but, unlike LLM-PFC, these model instances are not specialized to perform specific functions.
@@
  • Similar to LLM-PFC, reflexion (Shinn et al., 2023) involves an element of self evaluation of proposed policies, but this depends on interaction with the external environment to determine the outcome of each policy (whereas in LLM-PFC, this self evaluation process is entirely internal to the agent). The dependence on interaction with the external environment makes this approach unsuitable for the planning domain (planning is, by definition, performed internally).
@@
A.2 Experiment Details
@@
We implemented each of the modules using a separate GPT-4 (32K context, ‘2023-03-15-preview’ model index, Microsoft Azure openAI service) instance through a combination of prompting and few-shot in-context examples. We set Top-p to 0 and temperature to 0, except for the $\operatorname{Actor}$ (as detailed in section A.6.2). The $\operatorname{Search}$ loop explored $B=2$ branches recursively for a depth $L=2$.
@@
For ToH, we used two randomly selected in-context examples of three-disk problems, and a description of the problem in the prompts for all the modules. For the graph traversal tasks, we used two in-context examples for all modules, except for the $\operatorname{Actor}$ and $\operatorname{Evaluator}$ in the Steppath task, where we used three in-context examples, one each for 2-, 3-, and 4-step paths. The prompt also described the specific task that was to be performed by each module (e.g., monitoring, task decomposition). For more details about the prompts and specific procedures used for each module, see Section A.6.
@@
For three-disk problems, we allowed a maximum of $T=10$ actions per problem, and evaluated on 24 out of 26 possible problems (leaving out the two problems that were used as in-context examples for the $\operatorname{Actor}$). We also evaluated on four-disk problems, for which we allowed a maximum of $T=20$ actions per problem. The same three-disk problems were used as in-context examples, meaning that the four-disk problems tested for out-of-distribution (OOD) generalization. For the graph traversal tasks, we allowed a maximum of $T=6$ actions per problem.
@@
We didn’t use a separate $\operatorname{Predictor}$ for the graph traversal tasks, since the action proposed by the $\operatorname{Actor}$ gives the next state. We also did not include the $\operatorname{TaskDecomposer}$ for these tasks, and did not use the $\operatorname{Search}$ loop for the Steppath task, as the model’s performance was already at ceiling without the use of these components.
A.3 Illustration of Planning Tasks
@@
A.4 Results Tables
Table 1: Results on Valuepath task.
@@ A.4 Results Tables
LLM-PFC 100 0 1.5 2 4.75
Table 2: Results on Steppath task.
@@
LLM-PFC 100 100 95 0 0 0 2.1 3.42 4.5
Table 3: Results on Detour task.
@@
GPT-4 ICL 38 3
LLM-PFC 92 0
Table 4: Results on Reward Revaluation task.
@@
GPT-4 Zero-shot 38 14
GPT-4 ICL 31 8
LLM-PFC 54 0
Table 5: Results on ToH. Note that we also include results here for the GPT-4 ICL baseline when prompted with 5 ICL examples (as opposed to 2 examples in the standard version of the baseline). Surprisingly, more ICL examples hurts performance on this task, perhaps due to overfitting to the specific examples (Hasanbeig et al., 2023).
@@
GPT-4 CoT + ICL 42 5 22 39
GPT-4 MAD 25 1 24 43
LLM-PFC 74 24 0 0
Table 6: Ablation study on ToH with 3 disks.
@@
LLM-PFC 74 0
w/o Task Decomposer 50 0
w/o Tree Search 32 0
w/o Monitor 27 31
A.5 Analysis of Failure Modes
Failures in LLM-PFC largely stem from failures in the $\operatorname{TaskDecomposer}$, $\operatorname{Actor}$, and $\operatorname{Evaluator}$ modules. The $\operatorname{TaskDecomposer}$ sometimes fails to identify effective subgoals that will ultimately move the agent toward the final goal; the $\operatorname{Actor}$ sometimes fails to propose effective moves that move the agent toward the goal or subgoal; and the $\operatorname{Evaluator}$ sometimes fails to assign values that result in the selection of the best possible move as proposed by the $\operatorname{Actor}$. This is in contrast to the $\operatorname{Monitor}$ (responsible for detection of invalid moves), $\operatorname{Orchestrator}$ (responsible for determining whether a goal or subgoal has been achieved), and $\operatorname{Predictor}$ (responsible for predicting the states that will result from a proposed action), which all perform nearly perfectly in all tasks.
A.6 Prompts and In-context Examples
A.6.1 Task Decomposer

### v4

@@ 6 Conclusion and future directions
In this work, we have proposed the MAP architecture, a modular agentic approach aimed at improving planning with LLMs. In experiments on four challenging domains, we found that MAP significantly improved multi-step planning and decision-making performance over other LLM methods (e.g., Chain of Thought, Multi-Agent Debate, Tree of Thought). While these results represent a significant step forward, there is still room for improvement. In particular, while improving performance significantly, the model still has less than optimal performance on Tower of Hanoi, the Reward Revaluation graph traversal task, and the PlanBench benchmark (Valmeekam et al., 2023) (see Section A.6 for discussion of failure modes). This may be due in part to the inherent limitations of prompting and in-context learning as methods for the specialization of MAP’s modules. A promising avenue for further improvement may be to jointly fine-tune smaller open-source LLMs to serve as modules across a range of diverse tasks, rather than relying only on black box methods (as with GPT-4). This approach would also eliminate the need for task-specific prompts, and may further improve zero-shot planning on novel tasks.
@@
An additional limitation of the current implementation is the computational cost incurred by the model (see Section A.8). Although this aligns well with the deliberative nature of controlled (i.e., ‘system 2’) processes (Kahneman, 2011), it would nevertheless be desirable to find ways to reduce these costs. In Section A.8, we present results from a version of MAP that achieves significantly improved efficiency, while retaining the same level of performance, by caching and re-using the outputs of some modules. We also found that MAP is effective when used with a smaller model (Llama3-70B), though performance was not as strong as the version that used GPT-4. Further improvements may result from fine-tuning smaller models to perform the specialized roles of each module.
Finally, it is interesting to consider the close parallels between our proposed approach and the neural basis of human planning and decision-making. In the human brain, planning is generally thought to depend on the prefrontal cortex (PFC) (Owen, 1997; Momennejad et al., 2018; Momennejad, 2020; Russin et al., 2020; Brunec & Momennejad, 2022; Mattar & Lengyel, 2022), a region in the frontal lobe that is notably most developed in humans and is broadly involved in executive function, decision-making, and reasoning (Miller & Cohen, 2001). Research in cognitive neuroscience has revealed the involvement of several subregions or modules within the PFC that appear to be specialized to perform certain functions, many of which are closely aligned with some of the modules in our proposed approach. These include the Anterior Cingulate Cortex, which is known to play a role in conflict monitoring (Botvinick et al., 1999), similar to our $\operatorname{Monitor}$ module; the Orbitofrontal Cortex, which plays a role in state prediction and state evaluation (Wallis, 2007; Schuck et al., 2016), similar to our $\operatorname{Predictor}$ and $\operatorname{Evaluator}$ modules; and the Anterior PFC, which plays a role in task decomposition and coordination, similar to our $\operatorname{TaskDecomposer}$ and $\operatorname{Orchestrator}$ modules. Human planning then emerges through the coordinated and recurrent interactions among these specialized subregions, and, similar to our approach, the algorithms implemented via these interactions are closely related to RL (O’doherty, 2004; Daw et al., 2005; Valentin et al., 2007; Takahashi et al., 2011; Silvetti et al., 2014; Brunec & Momennejad, 2022; Wang et al., 2018; Botvinick et al., 2019). An exciting direction for future work is to consider how the present approach might further contribute to understanding the brain basis of planning and decision-making.
@@ Appendix A Appendix
A.1 Supplementary Algorithms
Function $\operatorname{ProposeAction}$($x,y,B$):
       $\sigma\leftarrow$ false // Initialize validity
[…]
       while $\sigma$ is false do
             $A\leftarrow\operatorname{Actor}(x,y,E,B)$ // Sample B actions
             $\sigma,\epsilon\leftarrow\operatorname{Monitor}(x,A)$ // Determine validity and provide feedback
             $E\leftarrow E\cup\{\epsilon\}$ // Accumulate feedback
       end while
return $A$
Algorithm 2 Action proposal loop. $\operatorname{ProposeAction}$ takes a state $x$ and a goal $y$ and generates $B$ potential actions $A=a_{b=1}\dots a_{b=B}$. This is implemented via a loop, in which the $\operatorname{Actor}$ first proposes potential actions, and the $\operatorname{Monitor}$ then assesses those actions according to certain constraints (e.g., task rules), providing feedback if any of the actions are deemed to be invalid. This continues until the proposed actions are considered valid. See Sections A.7.2 and A.7.3 for more details.
Function $\operatorname{Search}$($l,L,B,x,y$):
       $V_{l}\leftarrow\{\}$ // Initialize value record
       $\tilde{X}_{l}\leftarrow\{\}$ // Initialize next-state record
       $A_{l}\leftarrow\operatorname{ProposeAction}(x,y,B)$ // Propose B actions
       for $b$ in $1\ldots B$ do
             $\tilde{X}_{l}\leftarrow\tilde{X}_{l}\cup\{\tilde{x}_{lb}\}$ // Update next-state record
             $\Omega\leftarrow\operatorname{Orchestrator}(\tilde{x}_{lb},y)$ // Terminate search if goal achieved
             if $l<L$ and $\Omega$ is false then
                   $V_{l}\leftarrow V_{l}\cup\{v_{l+1}\}$ // Update value record
            else
                   $V_{l}\leftarrow V_{l}\cup\{v_{lb}\}$ // Update value record
             end if
       end for
       $a_{l}\leftarrow A_{l\operatorname{argmax}(V_{l})}$ // Select action
       $\tilde{x}_{l}\leftarrow\tilde{X}_{l{\operatorname{argmax}(V_{l})}}$ // Predicted next-state
return $a_{l},\tilde{x}_{l},v_{l}$
A.2 Extended Related Work
@@
In this section, we consider in more detail how MAP relates to existing black-box and agentic LLM approaches:
@@
  • Similar to MAP, both scratchpad (Nye et al., 2021) and chain-of-thought (CoT) (Wei et al., 2022b) decompose a problem into intermediate computations. However, unlike MAP, neither scratchpad nor CoT factorize these intermediate computations into specialized modules.
@@
  • Tree-of-thought (ToT) (Yao et al., 2023) introduces some degree of factorization, but the factorization is not as extensive as in MAP. The ‘generator’ module in ToT carries out a combination of the functions carried out by both the $\operatorname{Actor}$ (action proposal) and the $\operatorname{Predictor}$ (prediction of the states that will result from these actions) in MAP. The ‘evaluator’ module in ToT carries out a combination of the functions carried out by both the $\operatorname{Monitor}$ (error detection) and the $\operatorname{Evaluator}$ (prediction of state value) in MAP. ToT does not contain any component that carries out the functions of the $\operatorname{TaskDecomposer}$ (subgoal proposal) and the $\operatorname{Orchestrator}$ (autonomously determining when a goal or subgoal has been achieved).
  • Multi-agent debate (i.e., Society of Mind) (Du et al., 2023) involves the interaction of multiple LLM instances; but, unlike MAP, these model instances are not specialized to perform specific functions. Similarly, the Large Language Model-based Actor-Critic (LLaMAC) approach (Zhang et al., 2023a) involves interaction between many LLM agents, and incorporates more specialization (there is a ’critic’ module that coordinates the decision-making process across many ’actors’), though there is less module specialization than there is in MAP.
@@
  • Similar to MAP, reflexion (Shinn et al., 2023) involves an element of self evaluation of proposed policies, but this depends on interaction with the external environment to determine the outcome of each policy (whereas in MAP, this self evaluation process is entirely internal to the agent). The dependence on interaction with the external environment makes this approach unsuitable for the planning domain (planning is, by definition, performed internally).
@@
  • Some recent work has aimed to combine external tool use with LLM planning agents, including Task Planning and Tool Usage (TPTU) (Ruan et al., 2023; Kong et al., 2023). Because this work evaluates on tasks involving external tool use, it is not directly comparable with our approach, but we plan to extend our approach to incorporate tool use in future. We expect that MAP’s modular approach will afford improved performance in this domain, relative to the more minimal planning approaches that have previously been employed (i.e., involving planning in a single LLM agent).
A.3 Experiment Details
@@
We implemented each of the modules using a separate GPT-4 (32K context, ‘2023-03-15-preview’ model index for ToH and cogeval tasks, and 128K context, ‘0125-preview‘ model index for strategyQA and planbench tasks from Microsoft Azure openAI service) instance through a combination of prompting and few-shot in-context examples. We set Top-p to 0 and temperature to 0, except for the $\operatorname{Actor}$ (as detailed in section A.7.2). The $\operatorname{Search}$ loop explored $B=2$ branches recursively for a depth $L=2$.
@@
For ToH, we used two randomly selected in-context examples of three-disk problems, and a description of the problem in the prompts for all the modules. For the graph traversal tasks, we used two in-context examples for all modules, except for the $\operatorname{Actor}$ and $\operatorname{Evaluator}$ in the Steppath task, where we used three in-context examples, one each for 2-, 3-, and 4-step paths. For strategyQA, we didn’t use any in-context examples. For the logistics task from Planbench, we used two incontext examples for all modules except for $\operatorname{Actor}$ which used three in-context examples. For the mystery blocksworld (deceptive) task from Planbench, we used two incontext examples for all modules except for $\operatorname{Actor}$ and $\operatorname{Predictor}$ which used three in-context examples. For both the tasks from Planbench, we extracted the goal from the initial state conditions, and the state and the goal was separately fed as input to the modules as required. The prompt also described the specific task that was to be performed by each module (e.g., monitoring, task decomposition). For more details about the prompts and specific procedures used for each module, see Section A.7.
@@
For three-disk problems, we allowed a maximum of $T=10$ actions per problem, and evaluated on 24 out of 26 possible problems (leaving out the two problems that were used as in-context examples for the $\operatorname{Actor}$). We also evaluated on four-disk problems, for which we allowed a maximum of $T=20$ actions per problem. The same three-disk problems were used as in-context examples, meaning that the four-disk problems tested for out-of-distribution (OOD) generalization. For the graph traversal tasks, we allowed a maximum of $T=6$ actions per problem. For strategyQA we allowed $T=1$ action per problem. For the Planbench tasks we allowed a maximum of $T=4$ + number of actions in the ground truth plan.
@@
We didn’t use a separate $\operatorname{Predictor}$ for the graph traversal tasks, since the action proposed by the $\operatorname{Actor}$ gives the next state. We also did not include the $\operatorname{TaskDecomposer}$ for these tasks, and did not use the $\operatorname{Search}$ loop for the Steppath task, as the model’s performance was already at ceiling without the use of these components. For strategyQA we didn’t use the $\operatorname{Evaluator}$ or the $\operatorname{Orchestrator}$. For the Planbench tasks we didn’t use tree search, and for mystery blocksworld task we didn’t use the $\operatorname{TaskDecomposer}$.
A.4 Additional Description of Planning Tasks
Graph Traversal. We focused on a particular type of graph (Figure 4) with community structure (Schapiro et al., 2013) previously found to be challenging for a wide variety of LLMs. The first task, Valuepath, involves finding the shortest path from a given room to the room with the largest reward, while avoiding the room that has a smaller reward. A smaller reward and a larger reward are located at two different positions in the graph. We fixed the two reward locations, and created 13 problems based on different starting locations. The second task, Steppath, involves finding the shortest path between a pair of nodes. We evaluated problems with an optimal shortest path of 2, 3, or 4 steps. We generated 20 problems for each of these conditions by sampling different starting and target locations.
@@ A.4 Additional Description of Planning Tasks
Tower of Hanoi. In the original version of the Tower of Hanoi (ToH) task, there are three pegs and a set of disks of different sizes. The disks are stacked in order of decreasing size on the leftmost peg. The goal is to move all disks to the rightmost peg, such that the disks are stacked in order of decreasing size. There are a couple of rules that determine which moves are considered valid. First, a disk can only be moved if it is at the top of its stack. Second, a disk can only be moved to the top of another stack if it is smaller than the disks in that stack (or if the peg is empty). More complex versions of the task can be created by using a larger number of disks.
@@
A.5 Supplementary Results
Figure 6: Graph traversal results. ‘invalid’ indicates percentage of moves that are invalid ($\downarrow$ better). GPT-4 Zero-shot, ICL, CoT, and MAD baselines are deterministic, and therefore a single run was performed on all problems. Note that MAP did not employ tree search on the Steppath task, and did not employ task decomposition on any of the graph traversal tasks. Without tree search, MAP’s performance is deterministic, and therefore only a single run was performed on the Steppath task, whereas we performed 5 runs with ToT. Gray error bars reflect 95% binomial confidence intervals (for models evaluated on a single run). Dots reflect values of 0%. For Valuepath, Detour, and Reward Revaluation we performed 10, 10, and 5 runs respectively with MAP and ToT, and present average performance $\pm$ the standard error of the mean (black error bars).
Table 4: Results on Valuepath task. Values within brackets indicate best performance.
@@ A.5 Supplementary Results
GPT-4 CoT 91 4
MAD 73 9
ToT 55(100) 0
MAP 100 0 1.5 2 4.75
Table 5: Results on Steppath task. Values within brackets indicate best performance.
@@
GPT-4 CoT 95 79 47 0 7 14
MAD 82 79 39 8 6 14
ToT 67(100) 67(89) 50(63) 0 0 0
MAP 100 100 95 0 0 0 2.1 3.42 4.5
Table 6: Results on Detour task. Values within brackets indicate best performance.
@@
GPT-4 ICL 38 4
GPT-4 CoT 69 6
MAD 38 9
ToT 33 (100) 0
MAP 85 (100) 1
Table 7: Results on Reward Revaluation task. Values within brackets indicate best performance.
@@
GPT-4 Zero-shot 31 15
GPT-4 ICL 54 6
GPT-4 CoT 54 13
MAD 46 14
ToT 36 (77) 0
MAP 48 (69) 0
Table 8: Results on ToH. Note that we also include results here for the GPT-4 ICL baseline when prompted with 5 ICL examples (as opposed to 2 examples in the standard version of the baseline). Surprisingly, more ICL examples hurts performance on this task, perhaps due to overfitting to the specific examples (Hasanbeig et al., 2023). Values within brackets indicate best performance.
@@
GPT-4 CoT 42 5 22 39
MAD 25 1 24 43
ToT-MAP 6 (25) 1(4) 4 5
ToT 6 (25) 0 (0) 14 11
MAP 74 (100) 24 0 0
Table 9: Ablation study on ToH with 3 disks. Values within brackets indicate best performance.
@@
MAP 74 (100) 0
w/o Task Decomposer 50 (67) 0
w/o Tree Search 32 (42) 0
w/o Monitor 27 (33) 31
Table 10: Results on ToH with 3 disks with a smaller LLM (Llama3-70B). MAP with Llama3-70B even outperforms the best GPT-4 baseline (GPT-4 ICL).
Model % solved problems % invalid actions
Llama3-70B Zero-shot 19.2 33.8
Llama3-70B ICL 12.5 41.4
Llama3-70B CoT 29.2 33.3
GPT-4 ICL 46 12
Llama3-70B MAP 50 2
Table 11: Results on Mystery Blocksworld domain of PlanBench using a zero-shot version of MAP (no in-context examples were provided for any module).
Model % solved problems
GPT-4 Zero-shot 0.2
GPT-4 ICL 7.8
MAP Zero-shot 8.2
Table 12: Results on ToH 3 disks with baselines also provided with the goal recursion strategy (provided to the $\operatorname{Decomposer}$ of MAP).
Model % solved problems
GPT-4 Zero-shot 11
GPT-4 Zero-shot w/ goal recursion 23
GPT-4 ICL 46
GPT-4 ICL w/ goal recursion 46
GPT-4 CoT 42
GPT-4 CoT w/ goal recursion 50
MAP w/ goal recursion 74
Table 13: Results on Planbench with ToT on a subset of problems. Values within brackets indicate best performance.
Model Logistics (30 problems) Mystery BW (100 problems)
ToT 10.4 (16.7) 0.6 (3)
MAP 53.3 35
Table 14: Results on ToH with 3 disks varying the tree search hyperparameters of MAP. Mean and standard error are reported across 10 runs for computational cost and performance.
Model Num. calls Num. input tokens Num. output tokens % solved
MAP ($B=2,L=1$) 32.04 $\pm$ 1.5 32,630.62 $\pm$ 1,318.4 2,197.37 $\pm$ 146.7 55 $\pm$ 6
MAP ($B=2,L=2$) 43.46 $\pm$ 2.4 49,537.37 $\pm$ 2,563.7 2,727.55 $\pm$ 214.2 72 $\pm$ 2
MAP ($B=2,L=3$) 65.65 $\pm$ 3.6 90,406.30 $\pm$ 4,784.7 3,845.88 $\pm$ 294.8 69 $\pm$ 2
A.6 Analysis of Failure Modes
To better understand the failure modes displayed by MAP, we analyzed the log files for the Tower of Hanoi (ToH) task (3-disk problems, run accuracy=75%, failed to solve 6/24 problems). We identified the following three general failure modes:
1. Incorrect decomposition: failure to identify subgoals that lie along the optimal path.
2. No progress: taking actions that either move away from or do not make progress toward a subgoal.
3. Falling into loops: visiting a state more than once.
These failure modes were not mutually exclusive. Table  15 shows the number of problems (out of 6 total failures) that involved these failure modes. We then identified which modules were responsible for these failures. The ‘incorrect decomposition’ failure mode was due to mistakes by the Decomposer (2/6 failures). The ‘no progress’ and ‘falling into loops’ failure modes were due to mistakes made both by the Actor and the Evaluator. Specifically, the Actor sometimes failed to propose at least one action that made progress toward the goal, and the Evaluator sometimes failed to select the action that made progress toward the goal. The Actor and Evaluator made these mistakes at least once for each of the 6 failures. Overall, we found that failures stemmed from errors made by the Decomposer, Actor, and Evaluator modules, whereas the other modules (Monitor, Predictor, and Orchestrator) performed perfectly.
Table 15: Results for number of failures (out of 6 total failures for a given run) of MAP on ToH with 3 disks for each of the three failure modes.
Incorrect decomposition No progress Falling into loops
2/6 6/6 5/6
A.7 Prompts and In-context Examples
A.7.1 Task Decomposer